OpenStudy (anonymous):

What is the recursive rule for the sequence: −22.7, −18.4, −14.1, −9.8, −5.5 …? (Points : 4) an = an + 1 - 4.3 an = an - 1 + 4.3 an = an - 1 - 4.3 an = an + 1 + 4.3

OpenStudy (ipwnbunnies):

Doesn't it look like a certain number is being subtracted from each term?

OpenStudy (ipwnbunnies):

I mean added. ._.

OpenStudy (anonymous):

yes but I kind find out see I was sick during time I was supposed to be studying this instead I was puking all day...

OpenStudy (ipwnbunnies):

Ok, look. -22.7 + 4.3 = -18.4 -18.4 + 4.3 = -14.1 You see where I'm going?

OpenStudy (anonymous):

yes I do .... so then the answer is The last one?

OpenStudy (ipwnbunnies):

Perhaps.

OpenStudy (ipwnbunnies):

I'm not sure what all that '+1' stuff is.

OpenStudy (ipwnbunnies):

Oh >.<

OpenStudy (anonymous):

yeah if I don't pass I fail and get sent to military school....

OpenStudy (ipwnbunnies):

Ugh, soz, we can't help with tests >.<

OpenStudy (anonymous):

its a practice test?

OpenStudy (ipwnbunnies):

Well, you already said it was a test. >->

OpenStudy (anonymous):

where?